Circular No. 5913

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN OPHIUCHUS B. A. Harmon, Marshall Space Flight Center, NASA; and W. S. Pa-

ciesas, University of Alabama at Huntsville, report for the Compton Observatory BATSE Team: “The hard x-ray transient GRO J1719–24 (= GRS 1716–249), which had been decreasing in intensity at ∼ 0.3 percent per day from a peak of 1.4 Crab (20-100 keV) on Sept. 30 (IAUC 5900), began dropping much more rapidly since ∼ Dec. 9. The intensities in the band 20-100 keV for Dec. 10, 13, and 15 were 1.1 (± 10 percent), 0.9 (± 15 percent), and 0.4 (± 20 percent) Crab, respectively. On Dec. 16-18, the source dropped below the BATSE 3σ 1-day detection limit of 0.1 Crab. Preliminary indications showed no marked change in spectral shape during the intensity decrease. Observations in other wavelength bands are strongly encouraged.”

OJ 287 AND 3C 279 E. I. Robson, Joint Astronomy Centre, Hilo; and J. A. Stevens report:

“James Clerk Maxwell Telescope monitoring of 3C 279 shows that the optical flare noted on IAUC 5911 appeared at 1.1 mm between Nov. 28 and Dec. 5. The flux rose by only 10 percent, with a further 10-percent rise by Dec. 23. UKIRT service observations of OJ 287 at 2.2 μm show an increase of more than a factor of 3 between 1993 January and November. JCMT data show a corresponding rise of about a factor of 2 at 0.8 mm and less at 1.1 mm, but between Nov. 21 and Dec. 23, OJ 287 remained steady at millimeter to submillimeter wavelengths.”

CORRIGENDA On IAUC 4983, ‘Comet Austin (1989c_{1})’, lines 7-10, the exponents for

r should all be negative values.

On IAUC 5666, ‘Nova Puppis 1991’, for Nov. 21.62, read Nov. 21.62 UT,

The galaxy UGC 1071 in which SN 1993ae lies (IAUC 5888) is also IC 126, a designation that should take precedence.

On IAUC 5901, ‘GRO J0422+32’, line 7, for [Nov.] 25.72, read [Nov.] 25.22,

On IAUC 5909, ‘OJ 287’, line 11, for 14.660, [15.4 read 15.660, [14.9

1993 December 27 Daniel W. E. Green